OpenStudy (lizz123):

FAN AND MEDAL!! WRITE IN SCIENTIFIC NOTATION!! (2 × 10 negative 5th power)(3 × 10 to the 4th power)

OpenStudy (anonymous):

0.00002 x 40000

OpenStudy (lizz123):

\[(2\times10^{-5})(3\times10^{4})\]

OpenStudy (lizz123):

@StudyGurl14

OpenStudy (anonymous):

lel

OpenStudy (studygurl14):

\(\large (2\times10^{-5})(3\times10^4)\rightarrow(2\times3)(10^{-5}\times10^4)\rightarrow the~answer\) Remember that.... \(\large x^m\times x^n=x^{m+n}\)

OpenStudy (lizz123):

So is the answer is: \[6\times10^{-9}\]

OpenStudy (lizz123):

@StudyGurl14

OpenStudy (studygurl14):

almost. But... \(\large -5+4\neq -9\)

OpenStudy (lizz123):

I am still confused

OpenStudy (lizz123):

@StudyGurl14

OpenStudy (lizz123):

@StudyGurl14

OpenStudy (studygurl14):

The answer is \(\large 6\times10^{-1}\) Because \(-5+4=-1\)

OpenStudy (lizz123):

ohh thnx @StudyGurl14